# FusionDTI: Fine-grained Binding Discovery with Token-level Fusion for Drug-Target Interaction

### v1

@@ Abstract
Predicting drug-target interaction (DTI) is critical in the drug discovery process. Despite remarkable advances in recent DTI models through the integration of representations from diverse drug and target encoders, such models often struggle to capture the fine-grained interactions between drugs and protein, i.e. the binding of specific drug atoms (or substructures) and key amino acids of proteins, which is crucial for understanding the binding mechanisms and optimising drug design. To address this issue, this paper introduces a novel model, called FusionDTI, which uses a token-level Fusion module to effectively learn fine-grained information for Drug-Target Interaction. In particular, our FusionDTI model uses the SELFIES representation of drugs to mitigate sequence fragment invalidation and incorporates the structure-aware (SA) vocabulary of target proteins to address the limitation of amino acid sequences in structural information, additionally leveraging pre-trained language models extensively trained on large-scale biomedical datasets as encoders to capture the complex information of drugs and targets. Experiments on three well-known benchmark datasets show that our proposed FusionDTI model achieves the best performance in DTI prediction compared with seven existing state-of-the-art baselines. Furthermore, our case study indicates that FusionDTI could highlight the potential binding sites, enhancing the explainability of the DTI prediction [^1] The complete code and datasets are available at: https://github.com/ZhaohanM/FusionDTI..
@@ 1 Introduction
Figure 1: A. An illustration of the FusionDTI model contains frozen encoders, the fusion module, and the classifier. The TF focuses on fine-grained interactions between tokens within and across sequences. B. This is a token-level interaction instance of HIV-1 protease and lopinavir. Lopinavir forms a hydrogen bond with residue Dd (Asp25) in the active site of the protease via its ketone molecule ([C][=O]). C. The attention map of TF visualises the weight between tokens, indicating the contribution of each drug atom and residue to the final prediction result.
@@
To address these challenges, we propose a novel model (called FusionDTI) with a Token-level Fusion (TF) module for an effective learning of fine-grained interactions between drugs and targets. In particular, our FusionDTI model utilises two pre-trained language models (PLMs), namely Saport [33] as the protein encoder that is able to integrate both residue tokens with structure token; and SELFormer [46] as the drug encoder to ensure that each drug is valid and contains structural information. To effectively learn fine-grained information from these contextual representations of drugs and targets, we explore two strategies for the TF module, i.e. Bilinear Attention Network (BAN) [18] and Cross Attention Network (CAN) [21, 37], to find the best approach for integrating the rich contextual embeddings derived from Saport and SELFormer. We conduct a comprehensive performance comparison against seven existing state-of-the-art DTI prediction models. The results show that our proposed model achieves about 6% accuracy improvement over the best baseline on the BinddingDB dataset. The main contributions of our study are as follows:
@@ 2.1 Drug-target Interaction Prediction
DTI prediction serves as an important step in the process of drug discovery [10]. Traditional biomedical measurements from wet experiments are reliable but have a notably high cost and time-consuming development cycle, preventing their application on large-scale data [49]. In contrast, identifying high-confidence DTI pairs by computational models markedly narrow down the search scope of drug candidate libraries, and aims to identify drugs most likely to bind to a target. Support vector machine (SVM) [9] and random forest (RF) [15] are two traditional computational models for DTI by concatenating fingerprint ECFP4 [29] and PSC features [6]. Later works focused on representation learning approaches, such as CNNs and GNNs [20, 25]. For example, DeepConv-DTI [20] employed CNNs and a global max-pooling layer to extract local protein sequence patterns. GraphDTA [25] used GNNs for drug graph encoding and CNNs for protein sequence encoding. More recently, MolTrans [16] introduced an adaptation of the transformer for encoding, further enhanced by a DIIN module [14] to learn fine-grained interactions. DrugBAN [3] incorporated a deep BAN [18] framework with domain adaptation to facilitate explicit pairwise fine-grained interaction learning between drugs and targets. In addition, BioT5 [26] has been proposed as a comprehensive pre-training framework that integrates cross-modelling in biology in the DTI task. Despite these advances, these models have not proposed an effective way to capture fine-grained interaction information in the DTI.
@@ 3.1 Model Architecture
Input: The initial inputs of drugs and targets are string-based representations. For protein $\mathcal{P}$, the SA vocabulary [33, 35] is employed, where each residue is replaced by one of 441 SA vocabularies that bind an amino acid to a 3D geometric feature to address the lack of structural information in the amino acid sequences. For drug $\mathcal{D}$, as mentioned in the previous section, we use the SELFIES, which is a formal syntax that always generates valid molecular graphs [19]. We provide the steps and code for obtaining SA and SELFIES sequences in Appendix 6.3.
@@ 3.2 Drug and Protein Encoders
PLMs have shown promising achievements in the biomedical domain leveraging transformers since they pay attention to contextual information and are pre-trained on large-scale biomedical databases. Therefore, we utilise Saport [33] as a protein encoder to encode protein input $\mathcal{P}$ of both the SA sequence and amino acid sequence. Meanwhile, SELFormer [46] is used as our drug encoder to encode the drug SELFIES input $\mathcal{D}$. Then these encoded protein representation $\mathbf{P}$ and drug representation $\mathbf{D}$ are further used as inputs for the later fusion module (Subsection 3.3). These rich contextual representations ensure that we can explore the fine-grained binding information effectively. To further justify this, we also compare our encoders with other existing protein language models (such as ESM-2b [22]) and molecular language models (such as MoLFormer [31] and ChemBERTa-2 [1]), and the results can be found in Section 4.7.
@@ 3.3.2 Cross Attention Network (CAN)
In summary, our CAN module combines multi-head, self-attention and cross-attention mechanisms to capture dependencies within individual sequences and between different sequences for a more nuanced understanding of interactions. In the results of Sections 4.3 and 4.5, we analyse and compare these two fusion strategies and different fusion scales in detail.
@@ 4.1 Datasets and Baselines
Three public DTI datasets, namely BindingDB [13], BioSNAP [48] and Human [24, 8], are used for evaluation, where each dataset is randomly split into training, validation and test sets with a 7:1:2 ratio. Since DTI is a binary classification task, we use AUROC (area under the receiver operating characteristic curve) [3, 16] and AUPRC (area under the precision-call curve) [20, 25] as the major metrics to evaluate a model’s performance.
We compare FusionDTI with seven baseline models in the DTI prediction task. These models include two traditional machine learning methods such as SVM [9] and Random Forest (RF) [15], as well as four deep learning methods including DeepConv-DTI [20], GraphDTA [25], MolTrans [16] and DrugBAN [3]. The latter four models employ the same two-stage process whereby the drug and target features are initially extracted by specialised encoders before being integrated for prediction. In addition, we also include the BioT5 [26] model, which is a biomedical pre-trained language model that could directly predict the DTI. Further details regarding the datasets, baseline models, and the methodology for generating drug SELFIES and protein SA sequences are provided in Appendix 6.3.
@@ 4.2 Effectiveness Evaluation for DTI Prediction
Table 1: Performance comparison of FusionDTI and the baselines on the BindingDB, Human and BioSNAP datasets. (Best, Second Best).
@@
We start by comparing our FusionDTI model (FusionDTI-CAN and FusionDTI-BAN) with seven existing state-of-the-art baselines for DTI prediction on three widely used datasets. Table 1 reports the comparative results. In general, our FusionDTI-CAN model performs the best on all metrics and all three datasets. A key highlight from these results is the exceptional performance of FusionDTI-CAN on the BindingDB dataset, where FusionDTI-CAN demonstrates superior metrics across the board: an AUROC of 0.989, an AUPRC of 0.990, and an accuracy of 96.1%. Note that the main difference between the FusionDTI-CAN model with others is the fusion strategy. Furthermore, although FusionDTI-BAN and DrugBAN have the same BAN module, FusionDTI-BAN performs better across all three datasets. These results highlight not only the marked enhancements of FusionDTI over other models on the BindingDB dataset but also its effectiveness in capturing fine-grained information on DTI. We consider the fine-grained interactions for each drug-target pair in the DTI prediction task, which is why FusionDTI uses the token-level fusion module. Our FusionDTI method is highly aligned with biomedical pathways the binding process relates to the specific atom or substructure interacting with the residue. Therefore, fine-grained interaction information effectively improves the performance of models in predicting DTI.
4.3 Comparison of the BAN and CAN Fusion Modules
There are two fusion strategies available: BAN and CAN, thus determining which one works better is a key step for establishing FusionDTI’s prediction effectiveness. We perform a fair comparison involving the same encoders, classifier and dataset. As shown in Figure 4, we compare BAN and CAN by employing two linear layers to adjust the feature dimensions of the drug and target representations. With the feature dimension increasing, the performance of FusionDTI-CAN continues to rise, while that of FusionDTI-BAN reaches a plateau. When the feature dimension is 512, both of the variants attain their peak positions with an AUC of 0.989 and 0.967, respectively. These results indicate that the CAN module seems to be better suited to the DTI prediction tasks and in capturing fine-grained interaction information. In contrast, BAN may not be able to fully capture fine-grained binding information between proteins and drugs, such as the specific interactions between the drug atoms and residues. Therefore, these findings suggest that the CAN strategy is more effective and adaptable to the complexities involved in DTI prediction, providing a superior performance, especially as the feature dimension scales.
@@ 4.5 Ablation Study
Figure 5: Performance evaluation of fusion scales.
Figure 6: Performance comparison of protein encoders.
Figure 7: Performance comparison of drug encoders.
@@ 4.6 Analysis of Fusion Scales
In assessing fusion representations, it is critical to determine whether a more fine-grained modelling enhances the predictive performance. Thus, we define a grouping function with the parameter g (Group size) for averaging per group tokens before the CAN fusion module. The g, representing the number of tokens per group, controls the granularity of the attention mechanism. Specifically, when g is set to 1, the fusion operates at the token level, where each token is considered independently. On the other hand, when g is set to 512, the fusion will run at the global level. We have the flexibility to control the fusion scale for the drug and protein representations, but this needs to meet the requirement that the token length is divisible by group size. As shown in Figure 7, as the number of tokens per group increases from 1 to 512 (Maximum Token Length), the FusionDTI model performance decreases accordingly. This also aligns with the biomedical rules governing drug-protein interactions, where the principal factor influencing the binding is the interplay between the key atoms or substructures in the drug and primary residues in the protein. In addition, the CAN module outperforms BAN consistently at various scale settings, indicating that CAN better access the information between the drug and target. Consequently, this supports that the more detailed the interaction information obtained between the drugs and targets by the fusion module, the more beneficial it is for the enhancement of the model’s prediction performance.
@@ 4.7 Evaluation of PLMs Encoding
The protein encoder and drug encoder are fundamental for the token-level fusion of representations, as these encoders are responsible for generating fine-grained representations to better explore interaction information. Our proposed model employs two PLMs encoding two biomedical entities: the drug and protein, respectively. In terms of the protein encoders, Figure 7 compares the the performance of the two protein encoders (Saprot [33] and ESM-2b [22]) in combination with three different drug encoders: ChemBERTa-2 [1], SELFormer [46] and MoLFormer [31]. From the figure, we find that Saprot consistently outperforms ESM-2b when combined with all three drug encoders. As can be seen in Figure 7, SELFormer achieves the best performance in encoding the drug sequences among the three advanced drug encoders. Notably, the top-performing combination is Saprot and SELFormer, hence our proposed FusionDTI uses them as drug and protein encoders.

### v2

@@ Abstract
Predicting drug-target interaction (DTI) is critical in the drug discovery process. Despite remarkable advances in recent DTI models through the integration of representations from diverse drug and target encoders, such models often struggle to capture the fine-grained interactions between drugs and protein, i.e. the binding of specific drug atoms (or substructures) and key amino acids of proteins, which is crucial for understanding the binding mechanisms and optimising drug design. To address this issue, this paper introduces a novel model, called FusionDTI, which uses a token-level Fusion module to effectively learn fine-grained information for Drug-Target Interaction. In particular, our FusionDTI model uses the SELFIES representation of drugs to mitigate sequence fragment invalidation and incorporates the structure-aware (SA) vocabulary of target proteins to address the limitation of amino acid sequences in structural information, additionally leveraging pre-trained language models extensively trained on large-scale biomedical datasets as encoders to capture the complex information of drugs and targets. Experiments on three well-known benchmark datasets show that our proposed FusionDTI model achieves the best performance in DTI prediction compared with eight existing state-of-the-art baselines. Furthermore, our case study indicates that FusionDTI could highlight the potential binding sites, enhancing the explainability of the DTI prediction.  [^1] The complete code, datasets and demo are available at: https://github.com/ZhaohanM/FusionDTI..
@@ 1 Introduction
Figure 1: A. An illustration of the FusionDTI model contains frozen encoders, the fusion module, and the classifier. The token-level fusion (TF) focuses on fine-grained interactions between tokens within and across sequences. B. This is a token-level interaction instance of HIV-1 protease and lopinavir. Lopinavir forms a hydrogen bond with residue Dd (Asp25) in the active site of the protease via its ketone molecule ([C][=O]). C. The attention map of TF visualises the weight between tokens, indicating the contribution of each drug atom and residue to the final prediction result.
@@
To address these challenges, we propose a novel model (called FusionDTI) with a Token-level Fusion (TF) module for an effective learning of fine-grained interactions between drugs and targets. In particular, our FusionDTI model utilises two pre-trained language models (PLMs), namely Saport [34] as the protein encoder that is able to integrate both residue tokens with structure token; and SELFormer [47] as the drug encoder to ensure that each drug is valid and contains structural information. To effectively learn fine-grained information from these contextual representations of drugs and targets, we explore two strategies for the TF module, i.e. Bilinear Attention Network (BAN) [19] and Cross Attention Network (CAN) [22, 38], to find the best approach for integrating the rich contextual embeddings derived from Saport and SELFormer. We conduct a comprehensive performance comparison against eight existing state-of-the-art DTI prediction models. The results show that our proposed model achieves about 6% accuracy improvement over the best baseline on the BindingDB dataset. The main contributions of our study are as follows:
@@ 2.1 Drug-target Interaction Prediction
DTI prediction serves as an important step in the process of drug discovery [10]. Traditional biomedical measurements from wet experiments are reliable but have a notably high cost and time-consuming development cycle, preventing their application on large-scale data [51]. In contrast, identifying high-confidence DTI pairs by computational models markedly narrow down the search scope of drug candidate libraries, and aims to identify drugs most likely to bind to a target. Support vector machine (SVM) [9] and random forest (RF) [16] are two traditional computational models for DTI by concatenating fingerprint ECFP4 [30] and PSC features [6]. Later works focused on representation learning approaches, such as CNNs and GNNs [21, 26]. For example, DeepConv-DTI [21] employed CNNs and a global max-pooling layer to extract local protein sequence patterns. Similarly, GraphDTA [26] used GNNs for drug graph encoding and CNNs for protein sequence encoding. Furthermore, MolTrans [17] introduced an adaptation of the transformer for encoding, further enhanced by a DIIN module [15] to learn fine-grained interactions. Additionally, DrugBAN [3] incorporated a deep BAN [19] framework with domain adaptation to facilitate explicit pairwise fine-grained interaction learning between drugs and targets. Moreover, BioT5 [27] has been proposed as a comprehensive pre-training framework that integrates cross-modelling in biology in the DTI task. More recently, SiamDTI [49] utilises a double-channel network structure to acquire local and global protein information for cross-field DTI prediction. Despite these advances, these models have not proposed an effective way to capture fine-grained interaction information in the DTI.
@@ 3.1 Model Architecture
Input: The initial inputs of drugs and targets are string-based representations. For protein $\mathcal{P}$, the SA vocabulary [34, 36] is employed, where each residue is replaced by one of 441 SA vocabularies that bind an amino acid to a 3D geometric feature to address the lack of structural information in amino acid sequences. For drug $\mathcal{D}$, as mentioned in the previous section, we use the SELFIES, which is a formal syntax that always generates valid molecular graphs [20]. We provide the steps and code for obtaining SA and SELFIES sequences in Appendix 6.3.
@@ 3.2 Drug and Protein Encoders
PLMs have shown promising achievements in the biomedical domain leveraging transformers since they pay attention to contextual information and are pre-trained on large-scale biomedical databases. Therefore, we utilise Saport [34] as a protein encoder to encode protein input $\mathcal{P}$ of both the SA sequence and amino acid sequence. Meanwhile, SELFormer [47] is used as our drug encoder to encode the drug SELFIES input $\mathcal{D}$. Then these encoded protein representation $\mathbf{P}$ and drug representation $\mathbf{D}$ are further used as inputs for the later fusion module (Subsection 3.3). These rich contextual representations ensure that we can explore the fine-grained binding information effectively. To further justify this, we also compare our encoders with other existing protein language models (such as ESM-2 [23]) and molecular language models (such as MoLFormer [32] and ChemBERTa-2 [1]), and the results can be found in Section 4.7.
@@ 3.3.2 Cross Attention Network (CAN)
In summary, our CAN module combines multi-head, self-attention and cross-attention mechanisms to capture dependencies within individual sequences and between different sequences for a more nuanced understanding of interactions. In the results of Sections 4.3 and 4.6, we analyse and compare these two fusion strategies and different fusion scales in detail.
@@ 4.1 Datasets and Baselines
Three public DTI datasets, namely BindingDB [14], BioSNAP [50] and Human [25, 8], are used for evaluation, where each dataset is split into training, validation, and test sets with a 7:1:2 ratio using two different splitting strategies: in-domain and cross-domain. For the in-domain split, the datasets are randomly divided. For the cross-domain setting, the datasets are split such that the drugs and targets in the test set do not overlap with those in the training set, making it a more challenging scenario where models must generalise to novel drug-target interactions. Since DTI is a binary classification task, we use AUROC (area under the receiver operating characteristic curve) [3, 17] and AUPRC (area under the precision-call curve) [21, 26] as the major metrics to evaluate a model’s performance.
We compare FusionDTI with eight baseline models in the DTI prediction task. These models include two traditional machine learning methods such as SVM [9] and Random Forest (RF) [16], as well as five deep learning methods including DeepConv-DTI [21], GraphDTA [26], MolTrans [17], DrugBAN [3] and SiamDTI [49]. The latter five models employ the same two-stage process whereby the drug and target features are initially extracted by specialised encoders before being integrated for prediction. In addition, we also include the BioT5 [27] model, which is a biomedical pre-trained language model that could directly predict the DTI. Further details regarding the datasets, baseline models, and the methodology for generating drug SELFIES and protein SA sequences are provided in Appendix 6.3.
@@ 4.2 Effectiveness Evaluation for DTI Prediction
Table 1: In-domain performance comparison of FusionDTI and the baselines on the BindingDB, Human and BioSNAP datasets (Best, Second Best).
@@
Table 2: Cross-domain performance comparison of FusionDTI and the baselines on the BindingDB, Human and BioSNAP datasets (Best, Second Best).
BindingDB Human BioSNAP
Method AUROC AUPRC Accuracy AUROC AUPRC AUROC AUPRC Accuracy
SVM .490$\pm$.015 .460$\pm$.001 .531$\pm$.009 .621$\pm$.036 .637$\pm$.009 .602$\pm$.005 .528$\pm$.005 .513$\pm$.011
RF .493$\pm$.021 .468$\pm$.023 .535$\pm$.012 .642$\pm$.011 .663$\pm$.050 .590$\pm$.015 .568$\pm$.018 .499$\pm$.004
GraphDTA .536$\pm$.015 .496$\pm$.029 .472$\pm$.009 .822$\pm$.009 .759$\pm$.006 .618$\pm$.005 .618$\pm$.008 .535$\pm$.024
DeepConv-DTI .527$\pm$.038 .499$\pm$.035 .490$\pm$.027 .761$\pm$.016 .628$\pm$.022 .645$\pm$.022 .642$\pm$.032 .558$\pm$.025
MolTrans .554$\pm$.024 .511$\pm$.025 .470$\pm$.004 .810$\pm$.021 .745$\pm$.034 .621$\pm$.015 .608$\pm$.022 .546$\pm$.032
DrugBAN .604$\pm$.027 .570$\pm$.047 .509$\pm$.021 .833$\pm$.020 .760$\pm$.031 .685$\pm$.044 .713$\pm$.041 .565$\pm$.056
SiamDTI .627$\pm$.027 .571$\pm$.024 .563$\pm$.033 .863$\pm$.019 .807$\pm$.040 .718$\pm$.055 .725$\pm$.054 .623$\pm$.070
BioT5 .651$\pm$.002 .653$\pm$.003 .621$\pm$.005 .856$\pm$.003 .853$\pm$.003 .720$\pm$.008 .718$\pm$.004 .715$\pm$.009
FusionDTI-BAN .659$\pm$.002 .663$\pm$.002 .633$\pm$.003 .784$\pm$.002 .790$\pm$.003 .723$\pm$.002 .721$\pm$.002 .756$\pm$.001
FusionDTI-CAN .675$\pm$.005 .676$\pm$.012 .649$\pm$.005 .801$\pm$.037 .803$\pm$.032 .748$\pm$.021 .766$\pm$.017 .734$\pm$.012
We start by comparing our FusionDTI model (FusionDTI-CAN and FusionDTI-BAN) with eight existing state-of-the-art baselines for DTI prediction on three widely used datasets. Table 1 reports the in-domain comparative results. In general, our FusionDTI-CAN model performs the best on all metrics across all three datasets. A key highlight from these results is the exceptional performance of FusionDTI-CAN on the BindingDB dataset, where FusionDTI-CAN demonstrates superior metrics across the board: an AUROC of 0.989, an AUPRC of 0.990, and an accuracy of 96.1%. Note that the main difference between the FusionDTI-CAN model and others is the fusion strategy. Furthermore, despite FusionDTI-BAN and DrugBAN both utilising the same BAN module, FusionDTI-BAN consistently outperforms DrugBAN on all datasets.
However, in-domain classification using random splits holds limited practical significance. Thus, we also evaluate the more challenging cross-domain DTI prediction, where the training data and the test data contain distinct drugs and targets. This setting precludes the use of known drug or target features when making predictions on the test data. As shown in Table 2, the performance of all models is diminished compared to the in-domain setting due to the reduced availability of information. Nevertheless, the FusionDTI-CAN model demonstrates outstanding performance in cross-domain DTI prediction on the BindingDB and BioSNAP datasets, highlighting its robustness in predicting novel drug-target interactions. For instance, on the BindingDB dataset, FusionDTI-CAN achieves the highest metrics with an AUROC of 0.675 and an AUPRC of 0.676. This underscores the effectiveness of the model’s fusion strategy in diverse and challenging scenarios. Similarly, despite sharing the BAN module, FusionDTI-BAN continues to outperform DrugBAN, further confirming the effectiveness of the FusionDTI framework in addressing cross-domain prediction challenges.
These findings highlight not only the substantial improvements of FusionDTI over existing approaches but also its effectiveness in capturing fine-grained information on DTI. The key to this success lies in FusionDTI’s token-level fusion module, which enables the model to consider fine-grained interactions for each drug-target pair. This fine-grained interaction information aligns closely with biomedical pathways, where binding events often depend on the specific atoms or substructures involved in interactions with residues. Therefore, the model’s ability to capture such fine-grained interactions significantly enhances its predictive performance for DTI.
4.3 Comparison of the BAN and CAN Fusion Modules
@@
There are two fusion strategies available: BAN and CAN, thus determining which one works better is a key step for establishing FusionDTI’s prediction effectiveness. We perform a fair comparison involving the same encoders, classifier and dataset. As shown in Figure 4, we compare BAN and CAN by employing two linear layers to adjust the feature dimensions of the drug and target representations. With the feature dimension increasing, the performance of FusionDTI-CAN continues to rise, while that of FusionDTI-BAN reaches a plateau. When the feature dimension is 512, both of the variants attain their peak positions with an AUC of 0.989 and 0.967, respectively. These results indicate that the CAN module seems to be better suited to the DTI prediction tasks and in capturing fine-grained interaction information. In contrast, BAN may not be able to fully capture fine-grained binding information between proteins and drugs, such as the specific interactions between the drug atoms and residues. Therefore, these findings suggest that the CAN strategy is more effective and adaptable to the complexities involved in DTI prediction, providing superior performance, especially as the feature dimension scales.
@@ 4.6 Analysis of Fusion Scales
In assessing fusion representations, it is critical to determine whether more fine-grained modelling enhances the predictive performance. Thus, we define a grouping function with the parameter g (Group size) for averaging tokens within each group before the CAN fusion module. The parameter g, representing the number of tokens per group, controls the granularity of the attention mechanism. Specifically, when g is set to 1, the fusion operates at the token level, where each token is considered independently. In contrast, when g is set to 512, the fusion occurs at a global level, considering the entire embedding as a single unit. We have the flexibility to control the fusion scale for the drug and protein representations, but the token length must be divisible by the group size. As shown in Figure 7, as the number of tokens per group increases from 1 to 512 (Maximum Token Length), the performance of the FusionDTI model declines accordingly. This also aligns with the biomedical rules governing drug-protein interactions, where the principal factor influencing the binding is the interplay between the key atoms or substructures in the drug and primary residues in the protein. Furthermore, the CAN module outperforms BAN consistently at various scale settings, indicating that CAN better accesses the information between the drug and target. Consequently, this supports that the more detailed the interaction information obtained between the drugs and targets by the fusion module, the more beneficial it is for the enhancement of the model’s prediction performance.
Figure 5: Performance evaluation of fusion scales.
Figure 6: Performance comparison of protein encoders.
Figure 7: Performance comparison of drug encoders.
@@ 4.7 Evaluation of PLMs Encoding
The protein encoder and drug encoder are fundamental for the token-level fusion of representations, as these encoders are responsible for generating fine-grained representations to better explore interaction information. Our proposed model employs two PLMs encoding two biomedical entities: the drug and protein, respectively. In terms of the protein encoders, Figure 7 compares the the performance of the two protein encoders (SaProt [34] and ESM-2 [23]) in combination with three different drug encoders: ChemBERTa-2 [1], SELFormer [47] and MoLFormer [32]. From the figure, we find that SaProt consistently outperforms ESM-2 when combined with all three drug encoders. As can be seen in Figure 7, SELFormer achieves the best performance in encoding the drug sequences among the three advanced drug encoders. Notably, the top-performing combination is SaProt and SELFormer, hence our proposed FusionDTI uses them as drug and protein encoders.